OpenStudy (anonymous):

Find f(6) in the equation (gof)(6).

OpenStudy (anonymous):

whats the equation?

OpenStudy (anonymous):

just (gof)(6)

OpenStudy (anonymous):

@Gerardo_cast23

OpenStudy (anonymous):

there has to be an equation that goes with it.. they have to tell us what is g(x) and f(x) for us to find the answer

OpenStudy (anonymous):

oh hold on!

OpenStudy (anonymous):

f(x)= 1/x-5 and g(x)=x^2+2

OpenStudy (anonymous):

@Gerardo_cast23 does that help?

OpenStudy (anonymous):

Haha yes it does!

OpenStudy (anonymous):

woohoo! :)

OpenStudy (anonymous):

g(f(6)) so what we do is well first what is f(6)? you tell me.

OpenStudy (anonymous):

umm f(6) so you replace an x with 6? i dont really know how to work these :(

OpenStudy (anonymous):

lol. Well what you do is find f(6) so for f(x) plug in 6 everytime you see an x

OpenStudy (anonymous):

okay so 1/6-5 and g(6)=6^2+2?

OpenStudy (anonymous):

hold youre horses lol. So you found thaht f(6) is 1/6-5... So this equals 1. Now you have g(1)

OpenStudy (anonymous):

okay so now how do i find g(f(6))?

OpenStudy (anonymous):

Its a two step problem.. So we found f(6)=1.. so... g(f(6))= g(1)... So now for every x ing plug in 1 and thats your final answer

OpenStudy (shamil98):

its basically (1/6-5)^2+2

OpenStudy (anonymous):

g(1)=1^2+2

OpenStudy (shamil98):

or (1/1)^2 + 2 =3

OpenStudy (shamil98):

(gof) = plug in f into g

OpenStudy (anonymous):

what numbers do i plug into g? like what is f?

OpenStudy (shamil98):

f(x)= 1/x-5 and g(x)=x^2+2 g(x)= (1/x-5)^2 + 2

OpenStudy (shamil98):

but since f is f(6) you plug 6 into f first. 1/6-5 = 1

OpenStudy (shamil98):

then you plug f into the equation for g.

OpenStudy (anonymous):

okay so g(1(6))?

OpenStudy (shamil98):

the equation for g is g(x)=x^2 +2 you plug f into the x of this equation

OpenStudy (anonymous):

so g(1)=1^2+2? sorry im so dumb lol :(

OpenStudy (shamil98):

not g(1) ... its fine plug the value of f(x) into the equation of g.

OpenStudy (shamil98):

but yes g = 3

OpenStudy (anonymous):

im so sorry lol im trying really hard to understand so since f(x) is 6 would i fill in the x's with 6?

OpenStudy (anonymous):

oh.. lol

OpenStudy (shamil98):

your question is if f(6) find g(x). right? then they give the equations for f(x) and g(x)

OpenStudy (shamil98):

so you solve for f first, then you plug in the result of f into g.

OpenStudy (anonymous):

well the exact questions are Substitute the value you found in part 1 into g(x) to find g(f(6))

OpenStudy (shamil98):

yes thats correct.

OpenStudy (anonymous):

okay so let me see...

OpenStudy (shamil98):

f(6) just means replace the values of x with that number for example f(x)=x+1 = 7.

OpenStudy (anonymous):

oh okay! so can you explain to me how to break down and find the answer for g(f(6))? i just cant figure it out

OpenStudy (shamil98):

with the same equation?

OpenStudy (anonymous):

im guessing so, all i know is that "Substitute the value you found in part 1 into g(x) to find g(f(6))" is Part 2 to the first question asked on this page

OpenStudy (shamil98):

Okay, so g(f(6)) just means you substitue the value from the equation of f(x) when x for the equation of f(x) is 6.

OpenStudy (shamil98):

Then you substitute the sum of that into the equation for g.

OpenStudy (shamil98):

lets say f(x)=x+1 and g(x) = x^2 +3 f(6) (7)^2 + 3 = 52. get it?

OpenStudy (shamil98):

for this equation you plug 6 into the equation for f(x) which is f(x) = 1/x-5

OpenStudy (anonymous):

well i see how you got 52 but for my problem how would i plug in f into g?

OpenStudy (anonymous):

i know f(6) = 1

OpenStudy (shamil98):

you solve f(x)= 1/x-5, since x for this equation is 6 you plug it in f(x) = 1/6-5 f(x) = 1 then you plug in the answer into the equation for g

OpenStudy (shamil98):

the equation for g(x)= x^2 +1 so you plug 1 into the x value resulting in 1^2 +1 = 3

OpenStudy (anonymous):

oh i see! would 3 be the answer?

OpenStudy (shamil98):

yes

OpenStudy (anonymous):

Looks like I was replaced lol. Oh well.I hope youre understanding everything and learning

OpenStudy (anonymous):

@Gerardo_cast23 of course not :p i just need all the help i can get :)

OpenStudy (anonymous):

so just to make sure g(f(6)) = 3? @shamil98

OpenStudy (shamil98):

yes

OpenStudy (anonymous):

wow thanks!!

OpenStudy (shamil98):

yeah , no prob.

OpenStudy (anonymous):

with the same equations, would the expression for g(x) be 1?

OpenStudy (shamil98):

? you mean g(1) ?

OpenStudy (anonymous):

well the Part 1 question is "Find the expression for g(x)."

OpenStudy (anonymous):

and above that is says g(x)=x^2+2, so would x^2+2 be the answer for what is the expression of g(x)?

OpenStudy (shamil98):

Is this from a word problem?

OpenStudy (anonymous):

sort of i guess. this is exactly what it says: 4. For f(x)=1/x-5 and g(x) = x^2+2, find: A. Find the expression for g(x).

OpenStudy (shamil98):

http://i.imgur.com/BNVhugq.png

OpenStudy (shamil98):

if its just the expression for g(x) then yes g(x) = x^2 +2 is your answer lol.

OpenStudy (anonymous):

thanks!

OpenStudy (shamil98):

no probs.